Two refrigeration systems, one using Refrigerant 717 (Ammonia) and the other using Refrigerant 12 (Dichlorodifluoromethane). The upper and lower temperatures for both systems are 30ºC and -20ºC. Find the following for each:

i)          Which system has the largest refrigerant effect           

   ii)         Which system has the largest COP                                 

            iii)        The corresponding Carnot Efficiency for each system
Answer #1

Condenser Temperature (T) = 30 °C = 303 K

Evaporator Temperature (TE)--20 °C = 253 K

Part A

Calculation of refrigerant effect (RE

Consider Refrigerant medium is ammonia

h_{1}=h_{g}\textup{ at -20}\: \: ^{o}\textup{C}=1437.23\textup{ kJ/Kg}

h4- hf at 30 °C314.769 kJ/Kg

\textup{RE}=\left ( h_{1}-h_{4} \right )

RE = 1437.23-314.769= 1122.461 kJ/Kg

Consider Refrigerant medium is R-12

h_{1}=h_{g}\textup{ at -20}\: \: ^{o}\textup{C}=342.682\textup{ kJ/Kg}

h4- hf at 30 °C228.540 kJ/Kg

\textup{RE}=\left ( h_{1}-h_{4} \right )

\textup{RE}=342.682-228.540=114.142\textup{ kJ/Kg}

Based on the above calculation RE of Ammonia is greater than RE of R-12

\textup{Part B -- calculation of COP}

Consider Refrigerant medium is ammonia

h_{1}=h_{g}\textup{ at -20}\: \: ^{o}\textup{C}=1437.23\textup{ kJ/Kg}

h4- hf at 30 °C314.769 kJ/Kg

h_{2}=h_{g}\textup{ at 30}\: \: ^{o}\textup{C}=1486.14\textup{ kJ/Kg}

\textup{COP}=\frac{h_{1}-h_{4}}{h_{2}-h_{1}}

\textup{COP}=\frac{1437.23-314.769}{1486.14-1437.23}=22.94

Consider Refrigerant medium is R-12

h_{1}=h_{g}\textup{ at -20}\: \: ^{o}\textup{C}=342.682\textup{ kJ/Kg}

h4- hf at 30 °C228.540 kJ/Kg

h_{2}=h_{g}\textup{ at 30}\: \: ^{o}\textup{C}=363.566\textup{ kJ/Kg}

\textup{COP}=\frac{h_{1}-h_{4}}{h_{2}-h_{1}}

\textup{COP}=\frac{342.682-228.540}{363.566-342.682}=5.46

Based on the above calculation COP of Ammonia is greater than COP of R-12

\textup{Part C-- Carnot efficiency}\left ( \eta \right )

Condenser Temperature (T) = 30 °C = 303 K

Evaporator Temperature (TE)--20 °C = 253 K

\textup{Carnot efficiency}\left ( \eta \right )=\frac{T_{C}-T_{E}}{T_{C}}

\textup{Carnot efficiency}\left ( \eta \right )=\frac{303-253}{303}=0.165=16.5 %

\textup{carnot efficiency is same for both cases}